eviant:

PLS HELP!!! Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. A(n) = 12 + (n – 1)(3)

eviant:

@Shadow @JustSaiyan @563blackghost

Vocaloid:

n is the term number to find the first term, plug in n = 1 to find the fourth term, plug in n = 4 to find the tenth term, plug in n = 10 into A(n) = 12 + (n – 1)(3)

eviant:

so, 3,24,27 is the correct answer?

Vocaloid:

first term: A(n) = 12 + (n – 1)(3) = 12 + (0)(3) = 12 not 3 fourth term: A(n) = 12 + (4 – 1)(3) = 12 + 9 = 21 tenth term: A(n) = 12 + (10 – 1)(3) = 39

eviant:

thankx